Hi I have a couple of problems that I would like to know how to solve.
The involve the sum of two integers with a difference. You have to find the tow integers.
I have x + y = the sum
What do I do next?
Okay Checkley, I solved that one. The problem was The sum of two numbers is 38. Their difference is 12. What are the two numbers.
I solved this by using the subtraction method after first setting up the problem.
x+y =38
x-y=12
Once I subtracted I got x = 26 and y = 12. Is that correct?

Answer by checkley71(8403)

IT APPEARS YOU NEGLECTED TO ADD THE X+X=2X AND SUBTRACTED 38-12=26 INSTEAD OF ADDING 12+38=50.
X+Y=38
X-Y=12 NOW ADD THESE 2 EQUATIONS.
-------------------------------------
2X=50
X=25 ANSWER.
25+Y=38
Y=38-25
Y=13 ANSWER.
PROOF.
25-13=12
12=12